Circular No. 6811

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/1997 L1 AND P/1997 T3 The IAU Small Bodies Names Committee, after more than 15 weeks

of deliberation, has decided that the comet announced in the third item on IAUC 6754 and discussed further on IAUC 6759 is to be named P/1997 T3 (Lagerkvist-Carsenty). At the same time, it was agreed that the comet announced on IAUC 6677, and for which a name was indicated on IAUC 6681, will be renamed C/1997 L1 (Zhu-Balam). Although the renaming of a comet is not something to be taken lightly, the Committee wishes to make the point that these two cases, in effect very similar to each other, are quite unlike others that have heretofore arisen. Although there are many examples of comets judged by their initial discoverers to be asteroidal, yet found by subsequent observers to be cometary, there was no reason to believe, from the apparent sky positions and motions available, that C/1997 L1 was different from the main-belt minor planets reported in the same batch of data, or that P/1997 T3 was not a Trojan of the type being surveyed at the time. It is therefore reasonable to name these two comets, not only for the people who drew attention to the objects themselves, but also for the observers who recognized the physical cometary characteristics of the objects and publicized this fact in a timely manner. In appreciating that modern CCD and computer technology renders it likely that similar situations will arise in the future, the Committee hastens to point out that, if the initial positional information shows an apparently asteroidal object to be dynamically unusual, possibly—but not necessarily—by the time the discovery has been widely publicized, an observer who then recognizes cometary characteristics has no claim for recognition in the comet’s name.

COMETS C/1997 H3, C/1997 P3 AND C/1998 A1 C. St. Cyr, Naval Research Laboratory, on behalf of the SOHO-LASCO

Consortium (cf. IAUC 6685), reports the discovery of three more comets in C3 coronagraphic data. Two were found by D. A. Biesecker in an automatic search of the archive, the third by St. Cyr in recent data. None of the comets had a tail, and peak magnitudes were ∼ 6.6, 7.5 and 6.1, respectively. Measurements by Biesecker and St. Cyr have been reduced by G. V. Williams and are published in detail on MPEC 1998-B11, -B12 and -B13, together with orbit solutions by B. G. Marsden.

1997/98 UT  α_{2000}  δ_{2000}
Apr. 25.536  2^{h}25.^{m}4  +10^{o}10′  C/1997 H3
Aug. 7.640  8 52.6  +14 16  C/1997 P3
Jan. 12.286  19 53.5  −24 30  C/1998 A1
1998 January 23  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden